### v1

# Noncollinear Magnetic Multipoles in Collinear Altermagnets

Abstract
Altermagnets host an array of magnetic multipoles, which are often visualized and studied in the reciprocal space. In the real space, the relative phase of the multipoles of the spin-density around atoms determines whether a system is an altermagnet or a conventional antiferromagnet. In this study, we approach these real space multipoles in altermagnets using a combination of first principles calculations and group theory. We show that even in collinear magnets, the local spin density is necessarily noncollinear due to spin-orbit coupling. Moreover, the noncollinear contributions often provide a more direct illustration of the magnetic multipolar character of altermagnetism than the collinear contribution, which is dominated by the dipolar term. Our first principles calculations also show that 32-poles, in addition to the octupoles, can be visible in spin-density of d-wave altermagnets, and they must be taken into account in discussions of the macroscopic response. Finally, we elucidate the interplay between magnetism and subtle crystal structural distortions in perovskite altermagnets, which provide a fertile playground for studying phase transitions between antiferromagnetic and altermagnetic phases.
Introduction.- Altermagnetism (AM) has recently been identified as a magnetic phase distinct from both ferromagnetism and antiferromagnetism (AF) [1, 2, 3, 3, 4, 5]. Altermagnetic order breaks time-reversal symmetry $\mathcal{T}$ by itself, but preserves the combination of $\mathcal{T}$ with some crystallographic symmetries, ensuring compensation of the net magnetic dipole moment. These symmetries are often non-symmorphic, and cannot be inversion or translation, unlike in AF. This causes otherwise spin-degenerate bands to split in most of the Brillouin zone even in the absence of spin-orbit coupling (SOC), while remaining degenerate on particular nodal lines and planes. Analogous to the nodal structure found in unconventional singlet superconductors, these spin-splitting patterns in altermagnets are labeled as d-wave, g-wave, etc., corresponding to angular momentum $l=2$, $l=4$, etc, which in turn determine the number of nodal planes in the spin-splitting throughout the Brillouin zone [6, 7, 8, 9].
This connection between nodal behavior and angular momentum motivated the identification of multipolar order parameters for altermagnetism [10, 11, 12, 13, 14]. On the experimental side, there have been efforts and proposals to probe magnetic multipoles through perturbations that assess matrix elements in observable tensors like piezomagnetism [15, 16], second-order magnetoelectric effect [17, 18, 19, 20] and the magneto-optic Kerr effect [21, 22]. On the theoretical side, magnetic multipoles have been studied in metals via effective momentum-space models in connection with Pomeranchuk instabilities [23, 24, 25], and in insulators via an atomic description of $f$ or $d$ electron levels subjected to crystalline fields and SOC [26, 27, 28]. However, visualization and calculation of AM multipolar order parameters from first principles in real materials in real space is rather rare [29, 30, 31], and their behavior under SOC has not been studied in detail.
In this Letter, we approach this problem using a combination of group theory and density functional theory (DFT) calculations. First, we explicitly demonstrate that typical collinear altermagnets such as MnF2 and CrSb host higher-order atomic multipoles in real space, including not only octupoles but dotriacontapoles (32-poles), which are visible in the spatially resolved noncollinear spin density (NSD). Inspired by recent proposals that the microscopic origin of AM in MnF2 is the ferro-octupolar order [29], and in $\alpha$-Fe2O3 the ferro-32-polar order [30], we present a method to discern atomic magnetic multipoles allowed by symmetry, predict their relative order and quantify their magnitude. While similar methods [32] have been used to, e.g. quantify atomic magnetoelectric multipoles in actinides [33] and monopoles in lithium transition metal phosphates [34], we visualize the spin density and consider the effect of SOC in collinear AMs. In particular, we elucidate the connection of SOC to the NSD, and the induced multipole components. The NSD, while smaller than the collinear component by more than an order of magnitude, exhibits a nodal structure with distinct higher-order multipolar behavior consistent with symmetry, providing the most direct illustration of the AM order. We also argue that these higher-order multipoles are relevant to the macroscopic response. We present results for prototypical AMs MnF2 and CrSb, as well as perovskite KMnF3, where structural degrees of freedom –the fluorine octahedral rotations– tune the magnetic state from simple AF to AM, thus highlighting the interplay of magnetism with crystallographic symmetries in AMs.
Collinear magnetic multipoles.- Magnetically ordered phases of crystals can be classified using magnetic space groups (MSGs), which also determine the magnetic site symmetries, i.e. the magnetic point group of each atomic site. Similar to the Wyckoff positions commonly used in crystallography, each atomic site can be assigned a magnetic Wyckoff position, which are tabulated for all MSGs [35, 36]. While the magnetic Wyckoff position tables focus on the allowed dipole moments, the magnetic site symmetries also constrain the spin density $\textbf{m}(\textbf{r})$ around an atomic site. Generally, $\textbf{m}(\textbf{r})$ can be expanded in tesseral (cubic) harmonics of order $\ell$, forming a magnetic multipole of order $\ell+1$ [37, 38]. Hence, the allowed magnetic multipoles centered on an atomic site are determined by the site symmetry as well. However, the sign and magnitude of multipoles are not determined by the site symmetry. Below, we obtain the spin density $\textbf{m}(\textbf{r})$ in real space from DFT, and calculate the magnitude of multipoles by projecting its components within atomic spheres onto different tesseral harmonics [39] (see, e.g. Ref. [40, 29, 31] for related approaches). These magnitudes are given in the atomic units of $a_{0}^{\ell}\mu_{B}$, where $a_{0}$ is the Bohr radius and $\mu_{B}$ the Bohr magneton. In the following, we use a shorthand notation and, for example, denote the octupole with a spin density pattern along $z$ that depends on position r as $m_{z}(\textbf{r})\sim xy$ as $d_{xy}m_{z}$’. The magnitude of this octupole is $\sim\int xy\cdot m_{z}(\textbf{r})\penalty 10000\ d\textbf{r}$, where the integral is over the atomic sphere (see SM for details [39]). This approach does not take into account all contributions to the total unit cell multipoles, but it allows predicting which atomic multipoles are in-phase (IP) vs. out-of-phase (OP) on different atoms within a unit cell. Thus, it circumvents the issue that unit cell multipoles do not have well-established gauge-invariant formulations due to complications in defining the position operator in periodic systems [41, 34].
Figure 1: (a) Crystal and magnetic structures of MnF2 and CrSb. The blue lines represent the screw axes, the red and green arrows represents the local $x$, $y$ coordinate axes, and the black arrows represent the spins. (b) Isosurfaces of collinear spin density ($m_{z}(\mathbf{r})$) around a magnetic ion obtained from DFT. (c) Same as (b), but for the noncollinear ($m_{x}(\mathbf{r})$ and $m_{y}(\mathbf{r})$) components, with a smaller isosurface value.
In all systems we consider, the atomic dipole moments $\int\textbf{m}(\textbf{r})d\textbf{r}$ are along the crystallographic $c$ axis, and the spin density along this direction ($m_{z}(\textbf{r})$) is orders of magnitude larger than other components. Thus, if one focuses only on the dipole moments at the magnetic sites, the system is collinear. We thus refer to magnetic multipoles induced by $m_{z}(\textbf{r})$ as collinear multipoles, and focus on them first. Rutile MnF2 has the same magnetic structure as was originally proposed for RuO2, with anti-parallel Mn spins oriented along the $c$ axis as shown in Fig. 1a, giving rise to an effective ferro-octupolar order [29]. The MSG is $P4_{2}^{\prime}/mnm^{\prime}$ (#136.499), and the order parameter transforms as the $m\Gamma_{2}^{+}$ space group irrep of the paramagnetic group $P4_{2}/mnm.1^{\prime}$ (#136.496). Mn ions occupy the magnetic Wyckoff site 2a with site symmetry $mm^{\prime}m^{\prime}$, which allows $sm_{z}=\int m_{z}(\mathbf{r})d\mathbf{r}$ atomic dipoles, as well as higher order magnetic multipoles such as the $d_{xy}m_{z}$ octupole, even though the spin density $m_{z}(\textbf{r})$ from DFT is almost spherical (Fig. 1b).
Site symmetry alone does not distinguish between antiferromagnets and altermagnets: if all atomic octupoles had alternating signs on neighboring Mn ions, they would have given rise to antiferromagnetism only. For MnF2 to be altermagnetic, at least one type of octupole on the two Mn ions in the unit cell have to be in-phase (in a global coordinate system) to add up to a macroscopic octupole moment [29].
In the Supplemental Material (SM) [39], we provide details of the irrep subduction analysis [42], which shows that the only multipoles allowed on Mn ions, apart from the $A_{g}^{-}$ ones, transform as the $B_{1g}^{-}$ irrep of the $mmm.1^{\prime}$ point group; i.e. $m\Gamma_{2}^{+}\downarrow mmm.1^{\prime}\rightarrow A_{g}^{-}\oplus B_{1g}^{-}$. As shown in Fig. 1a, in MnF2 there is a nonsymmorphic $4_{2}$ screw symmetry that interchanges the two Mn atoms in the unit cell, and as a result of this symmetry, the local coordinate axes of these Mn ions are rotated by 90 degrees with respect to each other. This means that the local multipoles on either Mn ion are rotated with respect to each other, and even though all $B_{1g}^{-}$ multipoles on the two Mn ions are out-of-phase in the local coordinate systems, they are in-phase in a global coordinate system if they change sign under 90 rotation around the $c$ axis ($C_{4z}$). This in turn is determined from the angular dependence of the multipoles; e.g. $d_{xy}m_{z}$ is odd under $C_{4z}$ because $m_{z}$ is even under it but the orbital part $d_{xy}$ changes sign. The allowed collinear multipoles grouped according to whether they are in-phase (IP) or out-of-phase (OP) in global coordinates are:
$(\rm{OP})\,m\Gamma_{2}^{+}\,[B^{-}_{1g}]$ $:sm_{z},d_{z^{2}}m_{z},\;g_{z^{4}}m_{z},\;g_{x^{2}y^{2}}m_{z},$
$(\rm{IP})\,m\Gamma_{2}^{+}\,[B^{-}_{1g}]$ $:d_{xy}m_{z},\;g_{xyz^{2}}m_{z}.$
Here, $gm_{z}$ are 32-poles with $m_{z}(\textbf{r})$ forming a $g$-orbital-like ($l=4$) nodal pattern. The collinear dipoles are OP and hence cancel each other, as is expected for a compensated magnet, together with all the other OP multipoles. On the other hand, the octupole $d_{xy}m_{z}$ and the 32-pole $g_{xyz^{2}}m_{z}$ are IP and do not cancel. This is in line with the observation that collinear compensating moments in the rutile structure yield a d-wave altermagnet, i.e. the lowest non-zero multipole in the unit cell is an octupole [29].
Unlike MnF2, the hexagonal metal CrSb with MSG $P6_{3}^{\prime}/m^{\prime}m^{\prime}c$ is a g-wave altermagnet [43, 44, 45]. Even though the site symmetry $\bar{3}m^{\prime}$ allows collinear octupoles on the Cr ions in its AM phase, all of them are out-of-phase in the global coordinate system, and the lowest order in-phase multipole is the $g_{yz(y^{2}-3x^{2})}m_{z}$ dotriacontapole:
$(\rm{OP})\,m\Gamma_{4}^{+}\,[A^{-}_{2g}]$ $:sm_{z},\;d_{z^{2}}m_{z},\;g_{z^{4}}m_{z},$
$(\rm{IP})\,m\Gamma_{4}^{+}\,[A^{-}_{2g}]$ $:g_{yz(y^{2}-3x^{2})}m_{z},$
In other words, both the collinear dipoles and the octupoles cancel, and the lowest order collinear magnetic multipole that remains uncompensated is a 32-pole. Again, this is expected since the two magnetic ions are swapped by either the $6_{3}$ screw, shown in Fig. 1b, or a $c$-glide. The only multipole that is IP is the one whose orbital part is odd under both the 6-fold rotation and the diagonal $\{110\}$ mirrors.
Figure 2: (a) The isosurfaces for different components of DFT spin densities of MnF2. Orange and blue colors correspond to positive and negative values, and the isosurface values used for collinear ($m_{z}$) and noncollinear ($m_{x}$ and $m_{y}$) are not equal. (b) Examples of decompositon onto the tesseral harmonics of the spin density in MnF2. Red and blue data correspond to the Mn atoms on the corners and the body-center of the unit cell. Results for the rest of the multipoles are reported in the SM [39].
Spin-orbit coupling.- Both MnF2 and CrSb are collinear magnets with atomic dipoles parallel to the $c$ axis. However, any magnetic material can have a noncollinear spin density (NSD) in the general equivalent positions $(x,\penalty 10000\ y,\penalty 10000\ z)$ in the unit cell, because these positions do not have any symmetries that prohibit a local spin density $\textbf{m}(\textbf{r})$ with an arbitrary orientation. This fact, often ignored because the NSD cancels out when integrated around an atomic site in collinear magnets ($sm_{x}=\int m_{x}(\textbf{r})d\textbf{r}=sm_{y}=0$), is easy to demonstrate in DFT by simply plotting isosurfaces of the NSD. In Fig. 1c, we show the isosurfaces of $m_{x}(\textbf{r})$ and $m_{y}(\textbf{r})$ spin densities around Mn and Cr ions in MnF2 and CrSb from DFT. Even though MnF2 is a d-wave AM, it exhibits 32-poles in its real space NSD that do not cancel when summed over the two atoms, as shown in 2a. This means that even though the primary order paramater is octupolar, higher-order in-plane multipoles are dominant and affect physical observables that depend on in-plane moments. In CrSb, while non-canceling high-order multipolar behavior is also evident (see 3a), the spin density pattern is more complicated because of the partial occupation of the Cr orbitals. This necessitates projecting the NSD onto tesseral harmonics to single out the contributions from different multipoles.
We now consider the effect of SOC on the NSD. While the NSD is allowed by MSG symmetries, it is not invariant under a spin-only rotation $R_{s}(\theta)$ by and arbitrary angle $\theta$ around the $c$ axis. As a result, when SOC is ignored and $R_{s}(\theta)$ becomes a symmetry of the system, the noncollinear multipoles become zero. In Figures 2b and 3b, we present the magnitude of magnetic multipoles in MnF2 and CrSb, calculated from first principles, as a function of the SOC strength. In these figures, $\lambda_{SOC}=0$ corresponds to SOC being turned off, and $\lambda_{SOC}=1$ corresponds to the real value of SOC. When there is no SOC, the NSDs $m_{x}$, $m_{y}$ go to zero. Under finite SOC, the collinear multipoles are modified quadratically in $\lambda_{SOC}$, and only by a small amount that is unnoticeable in the scale of the figures [39]. On the other hand, while the noncollinear dipoles always remain zero by symmetry, the symmetry-allowed noncollinear octupoles and 32-poles become non-zero linearly in $\lambda_{SOC}$. Whether these components on the two different ions in the unit cell are in-phase or out-of-phase is consistent with the group theory predictions discussed earlier.
In MnF2, the noncollinear multipoles on the Mn ions allowed by symmetry are
$(\rm{OP})\,m\Gamma_{2}^{+}\,[B^{-}_{1g}]:\;$ $d_{xz}m_{x}+d_{yz}m_{y},$
$g_{xz(x^{2}-3y^{2})}m_{x}+g_{yz(3x^{2}-y^{2})}m_{y},$
$(\rm{IP})\,m\Gamma_{2}^{+}\,[B^{-}_{1g}]:\;$ $d_{yz}m_{x}+d_{xz}m_{y},$
$g_{yz(y^{2}-3x^{2})}m_{x}+g_{xz(x^{2}-3y^{2})}m_{y}.$
As was the case for the collinear ones, the noncollinear multipoles that are IP are the ones that are odd under 4-fold rotation, which is confirmed by the DFT results. In the $g$-wave altermagnet CrSb, non-zero multipoles on the Cr ions are
$(\rm{OP})\,m\Gamma_{4}^{+}\,[A^{-}_{2g}]:\;$ $d_{yz}m_{x}+d_{xz}m_{y},\;g_{yz^{3}}m_{x}+g_{xz^{3}}m_{y},$
$(\rm{IP})\,m\Gamma_{4}^{+}\,[A^{-}_{2g}]:\;$ $d_{xy}m_{x}+d_{x^{2}-y^{2}}m_{y},$
$g_{xyz^{2}}m_{x}+g_{(x^{2}-y^{2})z^{2}}m_{y},$
$g_{xy(x^{2}-y^{2})}m_{x}+g_{x^{2}y^{2}}m_{y},$
where, surprisingly, the lowest order IP noncollinear multipoles that do not cancel are octupoles that are odd under the 60 rotation.
The linear dependence of the noncollinear multipoles on $\lambda_{SOC}$ can be explained by the form of the commutator of $\mathcal{H}_{SOC}\sim\lambda_{SOC}\textbf{L}\cdot\textbf{S}$ with the unperturbed Hamiltonian $\mathcal{H}_{0}$, in terms of the orbital and spin ladder operators $L_{\mp}$ and $S_{\mp}$. This commutator is off-diagonal both in spin and orbital spaces, so starting from a state $|\Psi_{0}\rangle$ that is an eigenstate of $S_{z}$, the first order change in the wavefunction only includes terms with spin opposite to $|\Psi_{0}\rangle$. Hence, there is an in-plane spin density at first order in $\lambda_{SOC}$. However, the change in the non-spin-flip part of the wavefunction is nonzero only in the second order in the perturbation, which explains the trend observed in the collinear multipoles above. The emergence of a spin density with a $g$-orbital-like pattern in a system with spin-polarized electrons on the $d$-shell can also be explained similarly. In the SM [39], we provide a simple toy model to illustrate these observations more explicitly.
Figure 3: (a) The isosurfaces for different components of DFT spin densities of CrSb. Orange and blue colors correspond to positive and negative values, and the isosurface values used for collinear ($m_{z}$) and noncollinear ($m_{x}$ and $m_{y}$) are not equal. (b) Examples of decompositon onto the tesseral harmonics of the spin density in CrSb. Red and blue data correspond to the Cr atoms on the edge and the body-center of the unit cell, which are equal for IP multipoles (e.g. $d_{xy}m_{z}$) and opposite for OP multipoles (e.g. $m_{z}$ and $d_{xz}m_{x}+d_{yz}m_{y}$). Results for the rest of the multipoles are reported in the SM [39].
Structural control of altermagnetism.- Finally, we discuss a third material example to elucidate how IP multipoles emerge in an antiferromagnetic-to-altermagnetic transition. The perovskite crystal structure [9, 46, 47] is an extremely flexible platform to realize antiferromagnetism and altermagnetism. In particular, the perovskite KMnF3 [46] provides a suitable playground to study the dependence of altermagnetism on structural degrees of freedom. This compound undergoes a structural transition at $\sim$88 K from cubic $Pm\bar{3}m$ to tetragonal $I4/mcm$ space group. G-type collinear magnetic order also condenses at either the same or at a close temperature [48, 49]. The magnetic order parameter transforms as the $mR_{5}^{-}$ irrep of the parent $Pm\bar{3}m.1^{\prime}$ space group. Being a zone-boundary order, it preserves anti-translation (translation followed by time reversal) symmetry and hence is an AF that does not give rise to no spin-splitting.
In the tetragonal phase, the fluorine octahedra are rotated by $\sim 6^{o}$ around the $c$ axis, and the unit cell is doubled. The oxygen octahedral rotation order parameter transforms as the $R_{5}^{-}$ irrep, and has the same wavevector as the magnetic order parameter. In other words, it breaks the translational symmetry the same way that the magnetic order parameter does, and hence anti-translation is no longer a symmetry when both structural and magnetic order parameters condense. This leads to octahedral rotation induced altermagnetism in KMnF3, where the altermagnetic splitting is proportional to both the magnetic order parameter and the structural order parameter. This is evident in the DFT results. In Fig. 4, we show that the $m_{z}(\textbf{r})$ isosurfaces around each Mn ion rotate with the octahedra, and even though the total dipole density does not change much, the lowest order IP multipole, the 32-pole $g_{xy(x^{2}-y^{2})}m_{z}$, is linearly proportional to the octahedral rotation angle (Fig. 4b). This 32-pole transforms as the $m\Gamma_{1}^{+}$ irrep of $Pm\bar{3}m.1^{\prime}$, and it is induced by a trilinear coupling in the free energy: $\mathcal{F}\sim R_{5}^{-}(a,0,0)\cdot mR_{5}^{-}(a,0,0)\cdot m\Gamma_{4}^{+}$, consistent with its linear dependence on $R_{5}^{-}$ rotations [39].
We note that RbMnF3 has a Néel temperature that is comparable to that of KMnF3, but it retains the cubic crystal structure at all temperatures [50]. As a result, the partially Rb-substituted compound K1-xRbxMnF3 would likely have a lower temperature structural transition that is well separated from the magnetic one, and hence could serve as a model system to study antiferromagnetic to altermagnetic phase transitions.
Tensor observables.- The symmetry of response tensors is often described using the Jahn symbol [51]. The multipoles can also be connected to a Jahn symbol [14]: for example, the magnetic octupole $r_{i}r_{j}m_{k}$ is an axial and magnetic rank-3 object symmetric under exchange of $i$ and $j$, so it is denoted by the symbol $ae[V^{2}]V$. Thus, systems with a net magnetic octupole must allow non-zero elements of $ae[V^{2}]V$ response tensors. Examples of such tensors include piezomagnetism [52, 53] and the second-order magnetoelectric effect [54, 55].
The IP multipoles in MnF2 include the collinear component $d_{xy}m_{z}$, which corresponds to the piezomagnetic tensor element $\Lambda_{xyz}$. Since it is induced by the large collinear spin components, this element remains finite even in the absence of SOC, but it is strongly suppressed in this system by Luttinger compensation, which enforces equal numbers of spin up and down electrons at zero temperature as long as the gap remains open. On the other hand, since CrSb is d-wave only in the noncollinear components, strain can induce a linear magnetization only along the $x$ and $y$ directions, and the magnitude of this effect must be proportional to $\lambda_{SOC}$. The lowest rank tensors that are linearly proportional to the collinear 32-pole $g_{yz(y^{2}-3x^{2})}m_{z}$ are rank-5 tensors with $ae[V^{2}][V^{2}]V$. These include second-order piezomagnetism (magnetization proportional to the product of two different components of strain, e.g. $\varepsilon_{yz}$ and $\varepsilon_{xx}$), and strain derivative of the second order magnetoelectric effect.
Figure 4: (a) The $m_{z}(\mathbf{r})$ spin density around the Mn ions in KMnF3 is aligned with the F octahedra, and rotates without much distortion along with the octahedra. F ions are shown as grey spheres. (b) While the integrated $sm_{z}$ dipole density remains almost constant as a function of the octahedral rotation angle, the IP 32-poles $g_{xy(x^{2}-y^{2})}m_{z}$ turn on linearly with octahedral rotations. All other multipole components are reported in the SM [39].
Conclusions.- Using first principles calculations, we calculated the higher order atomic multipoles in altermagnetic materials by projecting the spin density onto multipoles, which provides a simple yet robust approach to calculate these multipoles in real materials and beyond simple models. Interestingly, even in systems with seemingly spherical spin density in the collinear direction, we found that the noncollinear components provide a more elucidative illustration of the presence and relative phase of altermagnetic multipoles. Our results also showed that 32-poles, which are often ignored in the discussion of $d$-wave altermagnets, can be in fact be the dominant contribution to spin density and in certain macroscopic properties. In KMnF3, the interplay between the 32-poles and the crystal structure adds altermagnetism to the list of phenomena induced by octahedral rotations in perovskites. In summary, our work underlines that the NSD, higher order multipoles, and spin-orbit coupling are all important features of altermagnets that need to be taken into account in real materials especially when predicting their macroscopic response.
Acknowledgments.- The authors acknowledge useful discussions with Libor Šmejkal and Hana Schiff. L.B. and T.B were supported by the NSF CAREER grant DMR-2046020. L.B. and T.B. were also partially supported by the National Science Foundation through the University of Minnesota MRSEC under Award Number DMR-2011401. R.M.F. was supported by the Air Force Office of Scientific Research under Award No. FA9550-21-1-0423.